OpenStudy (anonymous):

Given the function f(x) = 4(x+3) − 5, solve for the inverse function when x = 3. −21 −16 −10 −1

hartnn (hartnn):

do you know how to find inverse function ?

OpenStudy (anonymous):

hartnn that is y i asked the question cause i forgot how to find the inverse

hartnn (hartnn):

ok, let y = 4(x+3) − 5 isolate 'x' from this equation. start by adding 5 on both sides.

OpenStudy (anonymous):

ok so it would be 5+y=4(x+3)

hartnn (hartnn):

yes, now divide by 4 on both sides...

OpenStudy (anonymous):

ok so it would be \[\frac{ 5+y }{ 4 }=\frac{ x+3 }{ 4 }\]

hartnn (hartnn):

on tight side, 4(x+3)/4 is just x+3 got this ? the 4 gets cancelled

OpenStudy (anonymous):

oh ok

hartnn (hartnn):

so, (5+y)/4 = x+3 just subtract 3 from both sides and you would have isolated 'x' ! :)

OpenStudy (anonymous):

0k so when i subtract 3 do i subtract it from 5 or put it outside the perenthisis

hartnn (hartnn):

outside paranthesis, as there is a fraction, keep the -3 alone on left side

OpenStudy (anonymous):

ok so -3(5+y)/4=x

hartnn (hartnn):

did you multiply ? its actually (5+y)/4 -3 = x ok ? so your inverse function is just (5+y)/4 -3

hartnn (hartnn):

i mean (5+x)/4-3

OpenStudy (anonymous):

now how do i get it to one of my answer options

hartnn (hartnn):

we wanted inverse function when x=3 so just plug in x = 3 there!

OpenStudy (anonymous):

ok so it would be 8/4=2-3=-1

hartnn (hartnn):

correct! good work :)

hartnn (hartnn):

another method to do the same thing is to put f(x) = 3 4(x+3) − 5 = 3 will also give you x = -1 :)

OpenStudy (anonymous):

ok thanks man

hartnn (hartnn):

welcome ^_^